Project

General

Profile

NDOS Commissioning Meeting 2011/11/10

Attendance: Jon, Leon, Jan, Nick, Mat, Art, Martin, Kurt, Andrew, Gary

Agenda:

  1. Beam and Detector Update
  2. Installation Plans for the next week
  3. DAQ Update and Plans for next week
  4. DCS Update and Plans for next week
  5. AOB

Beam and Detector Report

  • Mat: beam was off most of day yesterday for MI work, but otherwise beam was fine
  • Art: MINOS ND had problems, so no shifters were in CR overnight, thus beam was
  • Jon: any lost APDs? Mat: an additional uncoated, sealed & cooled APD was lost this morning. Leon would like leaky coated APD and other failed units to be removed from detector. Should take < 30 minutes. Jon: anytime is fine. Kurt: new dispatcher to be rolled out;
  • Mat: nearline metrics have not yet been changed, and won't be until next week. Mat will leave a note on the whiteboard for shifters.

Installation plans for next week

  • No plans for next week.
  • Mat: we received 10 coated APDs from Hamamatsu, but coating thickness is being measured out at SciDet and these won't be available until ~Thanksgiving.
  • Gary: what quality are these? Are these rejects? Leon: yes, but in general they look pretty good (good enough for our needs now)

DAQ Update and Plans for next week

  • no plans reported

DCS Update and Plans for next week

  • PS at NDOS will be replaced next Tuesday morning. PS at teststand will be replaced on Wednesday. Work should take < 2 hours on detector.

AOB

  • Jan: entry by Mat in ECL: number of active channels was decreasing over time during a run, and problem went away after hitting the red button. Jon: in the future we should stop the run, instead of doing this in the middle of a run.